InfocomGameInformation PaulD.Smith paulsmith@dg.com ThetablebelowcontainsallInfocomgamesandallknowninformation abouteachgameandeachrevision.AlltheinformationbelowwascontributedbyPaulDavid Doherty<d.doherty@bamp.berlinet.in-berlin.de>,exceptmostoftheproductcodesanddifficulty ratingsandalltheinformationonLTOIIIfortheIBM-PC,whichwerecontributedbyDaniel Sharpe<dpsharpe@eos.ncsu.edu>,andtheinformationonLTOII&IIfortheMacintosh,which wascontributedbyme:-). Thecolumnsinthetableandtheirmeaningsareasfollows: GameTitle Specifiesthetitleofthegame.Gamesarelistedinalphabeticalorder. Aut (Author)Liststheinitialsoftheauthor(s)ofthegame.Authorsarelistedbelow,along withwhattheywereknowntobedoinginMarch,1992: DADouglasAdams-Workingonthe5thbookintheHHGttGseries DAnDaveAnderson-? BBBobBates-PresidentofLegendEntertainment MBeMichaelBerlyn-WritesSEGAgamesforAccolade MBMarcBlank-OwnsaComputerLandstoreinBend ABAmyBriggs-APhDcandidateinExperimentalPsychology JCJamesClavell-? SGStuGalley-WorksatThinkingMachinesCorp. JLJimLawrence-? DLDavidLebling-Writesaspreadsheet SMSteveMeretzky-WritesadventuresforLegendEntertainment BMBrianMoriarty-WorksatLucasfilmGames JONJeffO’Neill-? JWJerryWolper-? DL (DifficultyLevel):ManyoftheInfocomgames,especiallytheearlierones,wereratedwith adifficultylevel.Thelevelsare(inincreasingorderofdifficulty): IIntroductory SStandard AAdvanced EExpert -NoneGiven Notethattherearewidelydifferingopinionsastotheaccuracyoftheseratings,especially betweenE(Expert)andA(Advanced). Prodct (ProductCode):Theproductcodesforthegamesaregiveninthiscolumn.Theydenote thetypeofthegameandthenumberinthattype:thetypesare: @DATE@ Page1 InfocomGameInformation CSImmortalLegends(designedbyChallenge,Inc.) IAInteractive(Talesof)Adventure ICInteractiveComedy IEInteractiveEspionage IHInteractiveHorror IJInteractiveJourney(?) IKInteractiveJunior-level(Kids?) ILInteractiveLivingLiterature(Legend?) IMInteractiveMystery IRInteractiveRomance ISInteractiveScienceFiction IZInteractiveFantasy(ZorkSeries) RelNum (ReleaseNumber):Thereleasenumberofthegame.Allgameswiththesamerelease numberarecompatibleinthatsavedfilesfromonereleasenumbercanbeplayedbyother versionsofthatgamewiththesamereleasenumber. SerialNumber Theserialnumberofthegame:notetheserialnumberisactuallythedatetheversionwas released,inYYMMDDformat.Somegameshavemorethanoneserialnumberforthe samereleasenumber;thismeansthatsomeminorchangesweremadeandthegamewas re-released,butthatthetwoversionsarecompatible(seeabove). Z-CodeVersion SpecifieswhichversionofZIL(ZorkInterpretedLanguage)thegameiswrittenin.The versionsareasfollows: 3StandardSeries 4PlusSeries 5SolidGoldInteractiveFiction 6GraphicInteractiveFiction Iftheversionispostfixedwith(snd)itmeansthatversionofthatgamesupportssound. LTOINum. TheLostTreasuresofInfocompackageinwhichthegamewasre-released(ifany).Cur-rentlytherearetwopackages:IandII.TheyhavebothbeenreleasedfortheIBMPC, Macintosh,andAmigacomputersystems.Thetablecontainsinformationforallexcept theAmigaversionofLTOIII,sinceIhaven’treceivedinfofromanyonewhohasityet. Parenthesizedpostfixesspecifywhichversionofagameisincludedinthepackageforthe machine: aAmiga mMacintosh pIBMPC Nopostfixmeansthesameversionisavailableforallthreesystems. PI? Ifthegameisplayablewithversion3oftheportableInfocomdatafileinterpreterthen thiscolumncontainsa‘‘Y’’. @DATE@ Page2 InfocomGameInformationList GameTitleAutDProRelSerialZ-CodeLTOIPI? horLdctNumNumberVersionNum. 778508144II-798511224-- SMAIS5 AMindForeverVoyaging 548906066-- 748907146-- BB-IL1 Arthur:TheQuestForExcalibur JONSIM4978512183IY Ballyhoo 478709155-- 498709175-- 518709235-- 578712215I-BM-IZ8 BeyondZork MB-IE198710085II-BorderZone 868702124-- 1168706024II-DA-IC2 Bureaucracy MBe JW SIA3238408093IIY Cutthroats 268211083-Y 278310053IY MBEIM1 Deadline 108308103-Y 168311183I(a)Y 248511183-Y 298608203I(pm)Y SIZ4 Enchanter MB DL 478409143-Y 568412213-Y 588510023I(a)Y 598511083I(m)Y 318711195I(p)- SIS4 Hitchhiker’sGuideToTheGalaxy DA SM DAnSIA4378612153IIY HollywoodHijinx MBeAIA1228309163IY Infidel 308903226-- 838907066-- MB-IJ1 Journey 1188603253-Y 508607113-Y 598607303-Y 48804055-- SMSIC1 LeatherGoddessesofPhobos 2038705063I(pm)Y 2198709123(snd)I(a)Y 2218709183(snd)-Y DL-IH1 LurkingHorror SG 48609183-Y JL 98610223IY IIM5 Moonmist JON-IC3198707224II-NordandBertCouldn’tMakeHeadOrTailOf It 208307083-Y 298401183-Y 378510033I(pa)Y 108805315I(m)- SMSIS3 Planetfall AB-IR1268707303IIY PlunderedHearts 158405013-Y 158405223II(m)Y 168505153-Y 168506033II(p)Y I Seastalker SGJL IK1 IA2 218712145II(p)- 268801275(snd)II(m)- BB-CS1 Sherlock:TheRiddleOfTheCrownJewels @DATE@ Page3 InfocomGameInformationList GameTitleAutDProRelSerialZ-CodeLTOIPI? horLdctNumNumberVersionNum. JC 2958903216-- DL3228907066-- -IA5 Shogun 48401313-Y 68405083-Y 138510213I(a)Y 158511083I(pm)Y SMAIZ5 Sorcerer 638509163-Y 878609043IY DLEIZ6 Spellbreaker 158209013I(p)Y 178210213IY DLEIS1 Starcross SM-IS61078704303IY Stationfall DLAIM3148410053IY Suspect 58302223-Y 78304193-Y 88305213I(a)Y 88405213I(pm)Y MBeEIS2 Suspended 118605094-- 128609264II-BMSIZ7 Trinity 688505013II(m)Y 698509203II(p)Y 238807065-- BMIIZ0 Wishbringer-TheMagickStoneofDreams 138305243-Y 208311193-Y 218312083-Y 228409243IY SGSIM2 Witness 5?1-- 258205153(old)-Y 288210133-Y 308303303-Y 758309293-Y 768405093-Y 888407263IY 528711255-- SIZ1 ZorkI-TheGreatUndergroundEmpire MB DL 7UG3AU52-- 188205173-Y 228303313-Y 488409043IY AIZ2 ZorkII-TheWizardofFrobozz MB DL 108208183-Y 158303313-Y 158405183-Y 168232423-Y 178407273IY AIZ3 ZorkIII-TheDungeonMaster MB DL 2968810196I(m)- 3668903236I(a)- 3938907146I(p)- SM-IZ9 ZorkZero-TheRevengeofMegaboz @DATE@ Page4